# TPP-SD: Accelerating Transformer Point Process Sampling with Speculative Decoding

### v2

@@ A.3 Proof of Theorem 1
$\mathbb{P}(\hat{\tau}\in\mathcal{A}|\hat{\tau}\text{ is accepted})$ $=\frac{\mathbb{P}(\hat{\tau}\in\mathcal{A},\hat{\tau}\text{ is accepted})}{\mathbb{P}(\hat{\tau}\text{ is accepted})}=\frac{\mathbb{P}(\hat{\tau}\in\mathcal{A},\hat{\tau}\text{ is accepted})}{\mathbb{P}(\hat{\tau}\in\mathcal{T},\hat{\tau}\text{ is accepted})}.$ (18)
@@ C.1 Algorithm for TPP-SD
1 Inputs: Initial event $(t_{0},k_{0})$, end time $T$, number of draft events $\gamma$, target model $\mathcal{M}_{T}=\{\mathcal{E}_{T},g_{T}(\tau|\cdot),f_{T}(k|\cdot)\}$, draft model $\mathcal{M}_{D}=\{\mathcal{E}_{D},g_{D}(\tau|\cdot),f_{D}(k|\cdot)\}$;
@@ C.3.1 Sampling Synthetic TPPs
For the experiments in Section˜5.2, we evaluate sampling performance within the time window [0,100][0,100][ 0 , 100 ]. We report the discrepancy metrics $\Delta\mathcal{L}_{\text{ar}}^{\text{syn}}$ and $\Delta\mathcal{L}_{\text{sd}}^{\text{syn}}$, the KS statistics $D_{\text{KS}}$, the execution time $T_{\text{AR}}$ and $T_{\text{SD}}$, and the resulting speedup ratio $S_{\text{AR/SD}}$ as defined in Section˜5.1. We conducted experiments across three random seeds and report the mean for each metric to ensure result robustness.
@@ C.3.2 Sampling Real-world TPPs
For the experiments in Section˜5.3, we conduct evaluations within the same time window of [0,100][0,100][ 0 , 100 ]. In accordance with Section˜C.3.1, we report the discrepancy metric $\Delta\mathcal{L}^{\text{real}}$, the Wasserstein distance $D_{\text{WS}}^{t}$ and $D_{\text{WS}}^{k}$, the execution time $T_{\text{AR}}$ and $T_{\text{SD}}$ for AR sampling and TPP-SD, and the resulting speedup ratio $S_{\text{AR/SD}}$ defined in Section˜5.1. For computation of $D_{WS}^{t}$ and $D_{WS}^{k}$, we set $M=100,N=100$, which means that we fix the first $M=100$ events as history and perform $N=100$ independent repetitions of sampling the 101101101-th event, obtaining $\{(t_{i}^{\text{AR}},k_{i}^{\text{AR}}\}_{i=1}^{100}$ from AR sampling and $\{(t_{i}^{\text{SD}},k_{i}^{\text{SD}}\}_{i=1}^{100}$ from TPP-SD, and subsequently computing $D_{\text{WS}}^{t}$ and $D_{\text{WS}}^{k}$ as elaborated in Section˜5.1.
@@ C.4 Ablation Studies
[…]
@@
Table 4: Performance of TPP-SD with draft length $\gamma=10$ under different size of draft model. The distance metrics include KS statistic ($D_{\text{KS}}$) for synthetic datasets and Wasserstein Distance ($D_{\text{WS}}^{t}$, $D_{\text{WS}}^{k}$) for real datasets. We conducted experiments across three random seeds and report the mean for each metric. For all metrics, the best result is shown in bold, and the second best result is shown in underline.
@@ D.1 Why Not Use CIF-based Speculative Decoding?
D.2 Architectural Differences across Encoders
@@
where $[\mathbf{z}(t_{i})]_{j}$ is the $j$-th dimension of the temporal embedding vector $\mathbf{z}(t_{i})$, and $M,m$ are hyperparameters in AttNHP.
The attention mechanisms differ across encoder architectures. For clarity, we present the single-head attention formulation for each model. The $l$ history embedding of event $(t_{i},k_{i})$ is computed as:
@@
where $f(\cdot,\cdot)$ is a Gaussian kernel $f(\mathbf{q}^{(l)}(t_{j}),\mathbf{k}^{(l)}(t_{i}))=\exp(\frac{1}{\sqrt{D}}\mathbf{q}^{(l)}(t_{i})^{\top}\mathbf{k}^{(l)}(t_{j}))\in\mathbb{R}$. The key distinction lies in how query, key, and value vectors are derived at $l$. In THP and SAHP, $\mathbf{q}^{(l)}(t_{i})=\mathbf{Q}^{(l)}\mathbf{h}^{(l-1)},\mathbf{k}^{(l)}(t_{i})=\mathbf{K}^{(l)}\mathbf{h}^{(l-1)},\mathbf{v}^{(l)}(t_{i})=\mathbf{V}^{(l)}\mathbf{h}^{(l-1)}$ are obtained through linear projection, whereas in AttNHP:
@@
D.3 Differences between TPP Sampling and Long-horizon Prediction

### v3

@@ A.3 Proof of Theorem 1
[…]
@@ C.1 Algorithm for TPP-SD
1Inputs: Initial event $(t_{0},k_{0})$, end time $T$, number of draft events $\gamma$, target model $\mathcal{M}_{T}=\{\mathcal{E}_{T},g_{T}(\tau|\cdot),f_{T}(k|\cdot)\}$, draft model $\mathcal{M}_{D}=\{\mathcal{E}_{D},g_{D}(\tau|\cdot),f_{D}(k|\cdot)\}$;
@@ C.3.1 Sampling Synthetic TPPs
For the experiments in Section˜5.2, we evaluate sampling performance within the time window $[0,100]$. We report the discrepancy metrics $\Delta\mathcal{L}_{\text{ar}}^{\text{syn}}$ and $\Delta\mathcal{L}_{\text{sd}}^{\text{syn}}$, the KS statistics $D_{\text{KS}}$, the execution time $T_{\text{AR}}$ and $T_{\text{SD}}$, and the resulting speedup ratio $S_{\text{AR/SD}}$ as defined in Section˜5.1. We conducted experiments across three random seeds and report the mean for each metric to ensure result robustness.
@@ C.3.2 Sampling Real-world TPPs
For the experiments in Section˜5.3, we conduct evaluations within the same time window of $[0,100]$. In accordance with Section˜C.3.1, we report the discrepancy metric $\Delta\mathcal{L}^{\text{real}}$, the Wasserstein distance $D_{\text{WS}}^{t}$ and $D_{\text{WS}}^{k}$, the execution time $T_{\text{AR}}$ and $T_{\text{SD}}$ for AR sampling and TPP-SD, and the resulting speedup ratio $S_{\text{AR/SD}}$ defined in Section˜5.1. For computation of $D_{WS}^{t}$ and $D_{WS}^{k}$, we set $M=100,N=100$, which means that we fix the first $M=100$ events as history and perform $N=100$ independent repetitions of sampling the $101$-th event, obtaining $\{(t_{i}^{\text{AR}},k_{i}^{\text{AR}}\}_{i=1}^{100}$ from AR sampling and $\{(t_{i}^{\text{SD}},k_{i}^{\text{SD}}\}_{i=1}^{100}$ from TPP-SD, and subsequently computing $D_{\text{WS}}^{t}$ and $D_{\text{WS}}^{k}$ as elaborated in Section˜5.1.
@@ C.4 Ablation Studies
As illustrated in Figs.˜6(a) and 6(b), experiments with THP and SAHP encoders reveal patterns consistent with our findings in Section˜5.4: changes in draft length $\gamma$ minimally affect likelihood discrepancy $\Delta\mathcal{L}$ and distance metrics $D$, while acceptance rate $\alpha$ also diminishes with increasing $\gamma$, and speedup ratio $S_{\text{AR/SD}}$ exhibits a characteristic peak followed by a decline, eventually dropping below $1\times$ when draft lengths become excessive. Intermediate draft lengths ($\gamma\approx 5\text{--}15$) optimize computational efficiency across all experimental configurations.
@@
Table 4: Performance of TPP-SD with draft length $\gamma=10$ under different size of draft model. We use KS statistic ($D_{\text{KS}}$) for synthetic datasets and Wasserstein Distance ($D_{\text{WS}}^{t}$, $D_{\text{WS}}^{k}$) for real datasets. We conducted experiments across three random seeds and report the mean for each metric. For all metrics, the best result is shown in bold, and the second best result is shown in underline.
@@ D.1 Why Not Use CIF-based Speculative Decoding?
D.2 Necessity of Sampling from Large Target Models
A critical consideration in TPP sampling is the necessity of sampling from large draft models. For TPP data, large Transformer encoders (e.g., multi-layer, multi-head, with high-dimensional representations) are crucial for capturing complex temporal dependencies. Their larger parameter capacity allows them to model intricate data distributions and generate high-quality samples—albeit at the cost of slower autoregressive sampling. In contrast, smaller Transformer encoders (e.g., single-layer, single-head, with low-dimensional representations) are much faster at sampling due to reduced parameter capacity. However, they often struggle to model complex temporal patterns accurately, resulting in lower-quality samples that may not align well with the underlying data distribution.
To empirically demonstrate this performance gap, we evaluate a simple draft model $\mathcal{M}_{D}$ (single-layer, single-head THP) on both synthetic (Poisson, Hawkes, Multi-Hawkes) and real datasets (Taxi, StackOverflow, Amazon, Taobao), measuring sampling fidelity via $\Delta\mathcal{L}$ and $D_{KS}$, $D_{WS}$, and efficiency via wall-time.
As shown in Table˜5, the simplistic architecture of $\mathcal{M}_{D}$ struggles to capture temporal dependencies effectively even on synthetic datasets. Performance degrades further on real datasets, with significantly larger $\Delta\mathcal{L}$ and $D_{WS}$ values, as $\mathcal{M}_{D}$ cannot model the complex dependencies present in the data. Draft model alone performs poorly on this task, demonstrating the necessity of a large target model and the importance of our proposed TPP-SD framework, which leverages the speed of the draft model without sacrificing the fidelity provided by the target model.
Table 5: Evaluation of sampling methods across synthetic and real-world datasets. Autoregressive sampling with the draft model (AR (draft)) consistently yields lower sample fidelity (higher $\Delta\mathcal{L}$, $D_{KS}$, and $D_{WS}$) compared to the target model (AR (target)) and our proposed TPP-SD. For each metric, the best result is bolded, and the second-best is underlined.
Metric Sampling Type Poisson Hawkes Multi-Hawkes Taxi StackOverflow Amazon Taobao
$\Delta\mathcal{L}(\downarrow)$ AR (draft) 2.174 1.348 1.305 1.504 1.136 0.614 2.122
AR (target) 0.542 0.753 0.022 0.441 0.587 0.056 0.446
TPP-SD 0.349 0.276 0.321 0.065 0.231 0.129 0.033
$D_{KS}(\downarrow)$ AR (draft) 0.153 0.141 0.126 - - - -
AR (target) 0.038 0.044 0.069 - - - -
TPP-SD 0.036 0.043 0.053 - - - -
$D_{WS}^{t}(\downarrow)$ AR (draft) - - - 0.765 0.994 0.534 0.577
AR (target) - - - 0.201 0.470 0.189 0.236
TPP-SD - - - 0.082 0.391 0.078 0.076
$D_{WS}^{k}(\downarrow)$ AR (draft) - - - 0.768 0.821 0.759 1.384
AR (target) - - - 0.055 0.376 0.184 0.267
TPP-SD - - - 0.655 0.375 0.418 0.751
Wall-time $(\downarrow)$ AR (draft) 0.987 1.240 1.530 0.150 0.120 0.070 0.410
AR (target) 3.477 5.147 4.007 1.157 1.353 1.023 5.890
TPP-SD 1.647 2.547 1.893 0.453 0.700 0.290 3.460
D.3 Feasibility of Using Simple Poisson Draft
The standard thinning algorithm generates candidate events using a homogeneous Poisson process $\text{PoiP}(\overline{\lambda})$. This motivates a compelling question for our TPP-SD framework: is it possible to employ a draft model as simple as $\text{PoiP}(\overline{\lambda})$ to further accelerate sampling from complex target models without compromising fidelity?
We replace the 1-head-1-layer (abbreviated as 1-H-1-L) draft model to a simple homogeneous Poisson process draft model, whose parameter is estimated by maximum likelihood estimation on the data, i.e., $\hat{\lambda}=\frac{n}{T}$, where $n$ is the number of events and $T$ is the time span of the dataset. As for the experiment setup, we adopt THP as the backbone of target model for illustration simplicity. We conduct the additional experiments on 3 synthetic datasets (i.e., Poisson, Hawkes, and Multi-Hawkes) and 2 real-world datasets (i.e., Taobao and Amazon).
[…]
We can see from Table˜6 that, on synthetic datasets, TPP-SD with Poisson draft maintains high fidelity in sampling while achieving even more significant speedup across three datasets. The fast drafting process of Poisson compensates for the relatively lower acceptance rate. However, on real-world datasets, while TPP-SD with Poisson draft retains the sampling fidelity, we witness drop in speedup ratio. On complex real-world TPPs, the acceptance rate of Poisson draft is too low that it overwhelms the speedup from fast drafting. Therefore, in our setting of TPP-SD, we use a more robust drafting strategy, i.e., a slightly more complex draft model (1-layer-1-head THP) to ensure that TPP-SD can achieve high fidelity sampling across a wide range of datasets.
Table 6: Comparison between a 1-head-1-layer (1-H-1-L) Transformer draft model and a simple Poisson process draft within the TPP-SD framework. On synthetic data, the Poisson draft achieves a higher speedup at the cost of fidelity. On real-world data, its low acceptance rate leads to inferior performance in both speed and fidelity. For each metric, the best result between the two TPP-SD methods is bolded, and the second-best is underlined.
Dataset
Sampling
Method
$\Delta\mathcal{L}(\downarrow)$ $D_{KS}(\downarrow)$ $D_{WS}^{t}(\downarrow)$ $D_{WS}^{k}(\downarrow)$
Acceptance
Rate $\alpha(\uparrow)$
Speedup
$S_{AR/SD}(\uparrow)$
Poisson AR Sampling 0.542 0.038 - - - -
TPP-SD (1-H-1-L draft) 0.349 0.036 - - 0.740 2.110
TPP-SD (Poisson draft) 0.630 0.050 - - 0.370 3.230
Hawkes AR Sampling 0.753 0.044 - - - -
TPP-SD (1-H-1-L draft) 0.276 0.043 - - 0.720 2.113
TPP-SD (Poisson draft) 0.286 0.036 - - 0.520 5.430
Multi-Hawkes AR Sampling 0.022 0.069 - - - -
TPP-SD (1-H-1-L draft) 0.321 0.053 - - 0.750 2.117
TPP-SD (Poisson draft) 0.268 0.057 - - 0.390 3.610
Taobao AR Sampling 0.446 - 0.236 0.267 - -
TPP-SD (1-H-1-L draft) 0.033 - 0.076 0.751 0.260 1.597
TPP-SD (Poisson draft) 0.680 - 0.091 0.623 0.030 1.370
Amazon AR Sampling 0.056 - 0.189 0.184 - -
TPP-SD (1-H-1-L draft) 0.129 - 0.078 0.418 0.270 3.550
TPP-SD (Poisson draft) 0.066 - 0.090 0.086 0.010 2.230
D.4 Architectural Differences across Encoders
@@
where $\mathbf{z}(t_{i})]_{j}$ is the $j$-th dimension of the temporal embedding vector $\mathbf{z}(t_{i})$, and $M,m$ are hyperparameters in AttNHP.
The attention mechanisms differ across encoder architectures. For clarity, we present the single-head attention formulation for each model. The layer-$l$ history embedding of event $(t_{i},k_{i})$ is computed as:
@@
where $f(\cdot,\cdot)$ is a Gaussian kernel $f(\mathbf{q}^{(l)}(t_{j}),\mathbf{k}^{(l)}(t_{i}))=\exp(\frac{1}{\sqrt{D}}\mathbf{q}^{(l)}(t_{i})^{\top}\mathbf{k}^{(l)}(t_{j}))\in\mathbb{R}$. The key distinction lies in how query, key, and value vectors are derived at layer-$l$. In THP and SAHP, $\mathbf{q}^{(l)}(t_{i})=\mathbf{Q}^{(l)}\mathbf{h}^{(l-1)},\mathbf{k}^{(l)}(t_{i})=\mathbf{K}^{(l)}\mathbf{h}^{(l-1)},\mathbf{v}^{(l)}(t_{i})=\mathbf{V}^{(l)}\mathbf{h}^{(l-1)}$ are obtained through linear projection, whereas in AttNHP:
@@
D.5 Differences between TPP Sampling and Long-horizon Prediction